OpenStudy (anonymous):

The molar mass of HF is 20.01 g/mol. How many moles of Sn are required to react with 40 g of HF?

OpenStudy (abmon98):

Construct A balanced Equation Sn+2HF--->SnF2+H2

OpenStudy (abmon98):

Number of Moles=Mass(g)/Molar Mass(g/mol) Number of Moles of HF=40/20.01=1.99 moles Look at the Mole Ratio of HF to Sn

OpenStudy (abmon98):

The Mole Ratio is 2:1 so if we have 1.99 moles of HF then we must consume 1.99*1/2 moles of Sn